OpenStudy (anonymous):

solve (x+3)^3/2-11=53

OpenStudy (watchmath):

\((x+3)^3=64\) \((x+3)^3=4^3\) \(x+3=4\) \(x=1\)

OpenStudy (anonymous):

53 not 64

OpenStudy (watchmath):

after you move over the 11 you get 64

OpenStudy (anonymous):

okay

OpenStudy (watchmath):

ok but I made a mistake

OpenStudy (watchmath):

\((x+3)^{3/2}=64=4^3=16^{3/2}\) \(x+3=16\) \(x=13\)

OpenStudy (anonymous):

yeah because the first answer wasn't up there but that one is